# Testing strong-field QED with the avalanche precursor

### v1

@@ I Introduction
The interaction of charged particles and photons with strong electromagnetic (EM) fields has attracted considerable interest for almost a century.Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) Even before the modern formulation of quantum electrodynamics (QED), the process of pair production under the action of a strong EM field, vacuum nonlinearities, and the attainability of the maximum possible EM field strength were studied and discussed.Sauter (1931); Heisenberg and Euler (1936) At some point, it was understood that the processes of photon emission by charged particles and electron-positron pair production by photons in strong EM fields exhibit properties that were not encountered in perturbative QED, moreover, the interplay between these processes and particle dynamics in EM fields gives rise to new phenomena.Erber (1966); Baier and Katkov (1968); Ritus (1985) Before the invention of the laser, the studies of such processes were mostly believed to be of theoretical interest only. However, high-intensity lasers, which became available after the introduction of the chirped-pulse-amplification technique,Strickland and Mourou (1985) decisively changed the landscape of QED studies in the presence of strong EM fields, which are often referred to as Strong Field Quantum Electrodynamcis (SFQED). The flexibility of the laser EM field configuration and the ability of lasers to accelerate particles from underdense Esarey, Schroeder, and Leemans (2009) and solid density Daido, Nishiuchi, and Pirozhkov (2012) plasmas and then make them interact with strong EM fields of that very laser or a second one has paved the way for a plethora of experimental proposals to study QED in strong EM fields.
@@
Though the interaction of laser pulses with particle beams and plasma targets allows for a number of relevant experimental configurations, there are three characteristic setups that are discussed in the literature: laser-particle beam, laser-plasma, and laser-laser.Zhang et al. (2020); Gonoskov et al. (2022) Each of these setups is characterized by a unique interaction geometry and an interplay of SFQED effects with single-particle and collective dynamics. The main difference comes from the role of the laser. While in a laser-particle beam setup, the laser pulse serves as a target for the particle beam (which is in the spirit of studies at high-energy particle colliders), in the laser-plasma and laser-laser setups, the laser field serves not only as a target but also as an accelerator for charged particles. Furthermore, in the laser-particle beam setup, there are usually no plasma effects at play, since the particle densities in a beam do not reach high values. In contrast, the laser-plasma and laser-laser setups can be tailored to study collective effects. Up to now, only the laser-particle beam configurations was explored experimentally for moderate laser intensities, [^1] Here, we refer to the laser intensities being moderate as the ones $\sim 10^{20}$ W/cm2, whereas high intensities are usually $\sim 10^{22-24}$ W/cm2 as envisioned for the next generation multi-petawatt laser facilities. which means that mostly the photon production and electron beam depletion were detected, with occasional electron-positron pair production in collider-based experiments.Bula et al. (1996); Burke et al. (1997); Poder et al. (2018); Cole et al. (2018); Mirzaie et al. (2024); Matheron et al. (2024); Arrowsmith et al. (2024)
@@
Since SFQED processes such as photon emission and pair production depend strongly on the EM field strength, it was repeatedly shown in theoretical and computer simulation studies (see the reviewsDi Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) and references cited therein) that the utilization of high-power laser facilities Danson et al. (2019) will pave the way towards new unexplored regimes of particle interactions with strong EM fields. Novel multi-petawatt (PW) laser facilities such as ELI-NP,Radier et al. (2022); Tanaka et al. (2020) ELI Beamlines,Spinka et al. (2017); Weber et al. (2017) CoReLS,Sung et al. (2017) Apollon,Yao et al. (2025) SULF,Li et al. (2018) and the future NSF OPAL nsf can be used for efficient generation of ultra-bright high-energy radiation and particle beams by taking advantage of multiple photon emissions, prolific pair production, and re-acceleration of charged particles that have lost energy due to radiation emission (see e.g., Refs. Gonoskov et al. (2017); Magnusson et al. (2019a)).
@@
These interaction regimes can be revealed in QED cascades, which are arguably one of the most fascinating SFQED phenomena. A QED cascade is the process of an electron, positron, or photon energy transformation during their interaction with a strong EM field into multiple secondary particles (see Refs. Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) and references cited therein) and usually consists of a series of nonlinear Compton and Breit-Wheeler events following one after another in quick succession. The two basic cascade types are identified by the interaction mechanism: the shower and the avalanche (or self-sustained).
@@
A shower is a sequence of SFQED processes triggered by high-energy charged particles or photons that enter a strong field region.Bulanov et al. (2013); Blackburn and Murphy (2017); Mercuri-Baron et al. (2021); pou (2025) In such events, the strong field induces elementary processes, however, its influence on the particle motion is usually negligible. Thus, the initial particle energy is transformed into secondary particles and the total energy of all particles, initial and secondary, remains roughly constant during the interaction. In QED avalanches, charged particles (both initial and secondary) are continuously re-accelerated by the field.Bell and Kirk (2008); Kirk, Bell, and Arka (2009); Fedotov et al. (2010); Elkina et al. (2011) As a result, the total particle number and energy can grow exponentially over time, as the energy is drawn from the field. Both cascades operate in the so-called radiation-dominated (RD) regime of interaction. In this regime, the charged particle dynamics in a strong field are dominated by the emitted radiation, and a significant fraction of the field energy can be transferred into high-energy radiation.Bulanov et al. (2004); Di Piazza, Hatsagortsyan, and Keitel (2010); Bulanov et al. (2011a); Sokolov, Naumova, and Nees (2011); Ji et al. (2014); Liseykina, Peganov, and Popruzhenko (2025); Gonoskov et al. (2017)
@@
It is rather straightforward to relate the cascade types with the corresponding experimental setups: laser-particle beam interaction is suitable for showers,Pouyez et al. (2024) whereas avalanches can be generated in laser-laser interaction.Bashmakov et al. (2014); Gelfer et al. (2015); Grismayer et al. (2017); Mironov, Gelfer, and Fedotov (2021) It is worth noting that these associations are not set in stone and different types of cascades can occur either almost simultaneously or concurrently in different interaction geometries (e.g., see Refs. Mironov, Narozhny, and Fedotov (2014); Magnusson et al. (2019b); Samsonov, Kostyukov, and Nerush (2021)).
@@
Figure 1: Landscape of the interaction regimes for electrons injected in an electric antinode of a standing wave formed by two counterpropagating multi-petawatt laser pulses. The parameter space is spanned by the total field amplitude (in dimensionless units, $a_{*}=2a_{0}$) and the energy $\varepsilon_{e}=mc^{2}\gamma_{e}$ that an electron can gain when accelerated by the electric field of the standing wave. The top horizontal axis shows the corresponding values of the total laser power delivered by two circularly-polarized laser beams focused to the diffraction limit ($f/1$). The arrow at $P_{*}=50$ PW indicates the planned capabilities of NSF OPAL. The solid curves show $\varepsilon_{e}$ accounting for radiation reaction within the classical LL model (blue) and the SFQED prediction (magenta), given by Eqs. (2) and (4), respectively. The vertical lines highlight the transition to various interaction regimes. The star-shaped marker corresponds to the point beyond which prolific QED avalanches become accessible.
@@ II.1 Radiation-dominated regime in the classical picture
In the classical picture, the electron will be experiencing accelerating and decelerating phases during motion, so that its Lorentz factor will oscillate with the amplitude $\gamma_{e}\sim a_{*}$, where $a_{*}=2a_{0}$ and
@@
Here, $e>0$ is the elementary charge, $m$ is the electron mass, and $E_{0}$ and $\omega$ are the amplitude and frequency of a single laser beam. At $a_{*}>1$, an electron initially at rest will become relativistic within a subcycle of the field. As $a_{*}$ is increased further, the interaction changes qualitatively. Let us briefly discuss these interaction regimes, which we also illustrate in Fig. 1.
@@ II.2 Quantum radiation
[…]
@@
Within QED, it is natural to characterize the interaction regimes by the value of $\chi_{e}$. Emission becomes essentially quantum at $\chi_{e}\gtrsim 0.1$, [^2] The radiation intensity calculated within the classical and strong-field QED approaches overlaps at lower values of $\chi_{e}$.Ritus (1985); Baier and Katkov (1968); Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Bulanov et al. (2013); Niel et al. (2018) at which point the particle dynamics change. If the field is well below the Sauter-Schwinger limit ($E\ll E_{S}$) and the electron is ultrarelativistic, the electron motion can still be viewed as semiclassical. Namely, particles propagate along classical trajectories,Baier and Katkov (1968) but the RR is accounted for in random photon emission eventsElkina et al. (2011). These events can be treated as point-like within the locally constant field approximation,Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023) which we assume valid for the interaction regimes discussed in this work, as it generally requires $a_{0}\gg 1$.
@@
During the electron acceleration, $\chi_{e}$ varies and can increase with time, as well as $\gamma_{e}$. Hence, the emission probability rate is parametrized as $W_{\mathrm{nCE}}(t)=W_{\mathrm{nCE}}[\gamma_{e}(t),\chi_{e}(t)]$, and, more importantly, it can continuously grow along an electron’s semiclassical trajectory. As a result, the field energy can be efficiently converted into radiation. This is a striking feature of the standing-wave field configuration that defines the interaction. Notably, such restoration of $\chi_{e}$ is not revealed, for example, in the interaction of an electron with a single plane wave, as $\chi_{e}$ is conserved during semiclassic motion in between quantum emissions.Landau and Lifshitz The quantum parameter of each particle produced in an elementary QED event (either photon emission or pair production) is smaller than that of the particle initiating the event. These two points combined make the plane-wave setup (namely, a single laser), in principle, unsuited to access the important regimes investigated here.
[…]
@@ II.3 QED cascades
When $\chi_{\mathrm{em}}$ approaches 1 (marked by a star in Fig. 1), the probability of pair creation by the emitted photons becomes high. This signifies the transition from single-electron dynamics in the RD regime to the onset of QED avalanche-type cascades, which is essentially a many-body regime. As the initial and secondary charges continuously experience acceleration and emit hard photons, $e^{-}e^{+}$ pair yield can reach a steady exponential growth in time, $N\sim\exp(\Gamma t)$,Bell and Kirk (2008); Kirk, Bell, and Arka (2009); Fedotov et al. (2010); Mercuri-Baron et al. (2025) where $\Gamma$ is the so-called cascade growth rate. Estimates show that the field scale for reaching high-multiplicity cascades is of the order $a_{*}\sim\alpha a_{S}$edotov et al. (2010); Elkina et al. (2011) (shown in Fig. 1). However, it was demonstrated with simulationsRidgers et al. (2012, 2013); Mironov, Narozhny, and Fedotov (2014); Gelfer et al. (2015); Grismayer et al. (2017); Artemenko and Kostyukov (2017); Yu et al. (2018) and recent theoretical analysisMironov, Gelfer, and Fedotov (2021); Mercuri-Baron et al. (2025) that avalanche-type cascades can onset at significantly lower fields.
@@
The spatial distribution of the field, e.g., if laser beams are tightly focused, also affects the cascade growth rate $\Gamma$.Tamburini, Di Piazza, and Keitel (2017) As the interaction area for such fields is small, $e^{\pm}$-s and $\gamma$-s migrate from the focal area during the interaction at rates $\nu_{e,\gamma}\propto\omega^{-1}$.Jirka et al. (2016) This limits the pair yield in a cascade at lower fields,Tang (2024) and imposes a threshold condition for the steady state formation.Mercuri-Baron et al. (2025) Overall, $\Gamma$ depends on the field parameters parametrically via the rates of particle production and migration:Mercuri-Baron et al. (2025)
@@
$\Gamma\simeq\frac{4\pi}{\omega}\left[-1+\frac{1.52\pi^{2}\alpha a_{*}}{\omega\tau_{C}a_{S}}\exp\left(-1.92\frac{\alpha a_{S}}{a_{*}}\right)\right],$ (6)
@@
We emphasise that for efficient particle production, the cascade should reach a steady state. In short laser pulses, this is not guaranteed, especially if the field strength is close to $a_{\mathrm{th}}$. Nevertheless, $e^{-}e^{+}$ pair yield can be relatively high already at the onset stage. We denote this subregime as the avalanche precursor: a transition from the quantum RD regime to high-multiplicity avalanches. For the sake of an estimate, one may assume that it extends up to the field strength $a_{*}$ such that $\Gamma(a_{*})=2\pi/\omega$ (also shown in Fig. 1 as the vertical green dash-dotted line). The latter condition guarantees a high particle yield in one field cycle.

### v2

@@ I Introduction
The interaction of charged particles and photons with strong electromagnetic (EM) fields has attracted considerable interest for almost a century.Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) Even before the modern formulation of quantum electrodynamics (QED), the process of pair production under the action of a strong EM field, vacuum nonlinearities, and the attainability of the maximum possible EM field strength were studied and discussed.Sauter (1931); Heisenberg and Euler (1936) At some point, it was understood that the processes of photon emission by charged particles and electron-positron pair production by photons in strong EM fields exhibit properties that were not encountered in perturbative QED, moreover, the interplay between these processes and particle dynamics in EM fields gives rise to new phenomena.Schwinger (1951); Reiss (1962); Nikishov and Ritus (1964); Brown and Kibble (1964) The initial studies of strong-field effects were mainly of theoretical interest, motivated by astrophysical applications.Erber (1966); Baier and Katkov (1968); Ritus (1985) However, high-intensity lasers, which became available after the introduction of the chirped-pulse-amplification technique,Strickland and Mourou (1985) decisively changed the landscape of QED studies in the presence of strong EM fields, which are often referred to as Strong Field Quantum Electrodynamcis (SFQED). The flexibility of the laser EM field configuration and the ability of lasers to accelerate particles from underdense Esarey, Schroeder, and Leemans (2009) and solid density Daido, Nishiuchi, and Pirozhkov (2012) plasmas and then make them interact with strong EM fields of that very laser or a second one has paved the way for a plethora of experimental proposals to study QED in strong EM fields.
@@
Though the interaction of laser pulses with particle beams and plasma targets allows for a number of relevant experimental configurations, there are three characteristic setups that are discussed in the literature: laser-particle beam, laser-plasma, and laser-laser.Zhang et al. (2020); Gonoskov et al. (2022) Each of these setups is characterized by a unique interaction geometry and an interplay of SFQED effects with single-particle and collective dynamics. The main difference comes from the role of the laser. While in a laser-particle beam setup, the laser pulse serves as a target for the particle beam (which is in the spirit of studies at high-energy particle colliders), in the laser-plasma and laser-laser setups, the laser field serves not only as a target but also as an accelerator for charged particles. Efficient acceleration can be achieved in very different regimes and geometries,Gonoskov et al. (2022) e.g. in a wake created in laser-plasma interaction or directly by an electric field at the magnetic nodes of a standing wave formed in laser-laser head-on collision. The latter configuration is central in our work. In the laser-particle beam setup, there are usually no plasma effects at play, since the particle densities in a beam do not reach high values. In contrast, the laser-plasma and laser-laser setups can be tailored to study collective effects. Up to now, only the laser-particle beam configurations was explored experimentally for moderate laser intensities, [^1] Here, we refer to the laser intensities being moderate as the ones $\sim 10^{20}$ W/cm2, whereas high intensities are usually $\sim 10^{22-24}$ W/cm2 as envisioned for the next generation multi-petawatt laser facilities. which means that mostly the photon production and electron beam depletion were detected, with occasional electron-positron pair production in conventional accelerator-based experiments.Bula et al. (1996); Burke et al. (1997); Poder et al. (2018); Cole et al. (2018); Mirzaie et al. (2024); Matheron et al. (2024); Arrowsmith et al. (2024) New experiments within this paradigm are the E320 at SLACSalgado et al. (2021) and the LUXE at DESYcollaboration et al. (2024), which will explore previously unaccessible regimes of SFQED.
[…]
@@
Since SFQED processes such as photon emission and pair production depend strongly on the EM field strength, it was repeatedly shown in theoretical and computer simulation studies (see the reviewsDi Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) and references cited therein) that the utilization of high-power laser facilities Danson et al. (2019) will pave the way towards new unexplored regimes of particle interactions with strong EM fields. Novel multi-petawatt (PW) laser facilities such as ELI-NP,Radier et al. (2022); Tanaka et al. (2020) ELI Beamlines,Spinka et al. (2017); Weber et al. (2017) CoReLS,Sung et al. (2017) Apollon,Yao et al. (2025) SULF,Li et al. (2018) and the future NSF OPAL nsf (a) can be used for efficient generation of ultra-bright high-energy radiation and particle beams by taking advantage of multiple photon emissions, prolific pair production, and re-acceleration of charged particles that have lost energy due to radiation emission (see e.g., Refs. Gonoskov et al. (2017); Magnusson et al. (2019a)).
@@
These interaction regimes can be revealed in QED cascades, which are arguably one of the most fascinating SFQED phenomena. A QED cascade is the process of an electron, positron, or photon energy transformation during their interaction with a strong EM field into multiple secondary particles (see Refs. Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) and references cited therein) and usually consists of a series of nonlinear Compton and Breit-Wheeler events following one after another in quick succession. The two basic cascade types are identified by the interaction mechanism: the shower and the avalanche (or self-sustained).Mironov, Narozhny, and Fedotov (2014)
@@
A shower is a sequence of SFQED processes triggered by high-energy charged particles or photons that enter a strong field region.Bulanov et al. (2013); Blackburn and Murphy (2017); Mercuri-Baron et al. (2021); Pouyez et al. (2025a) In such events, the strong field induces elementary processes, however, its influence on the particle motion is usually negligible. Thus, the initial particle energy is transformed into secondary particles and the total energy of all particles, initial and secondary, remains roughly constant during the interaction. In QED avalanches, charged particles (both initial and secondary) are continuously re-accelerated by the field.Bell and Kirk (2008); Kirk, Bell, and Arka (2009); Fedotov et al. (2010); Elkina et al. (2011) As a result, the total particle number and energy can grow exponentially over time, as the energy is drawn from the field. Both cascades operate in the so-called radiation-dominated (RD) regime of interaction. In this regime, the charged particle dynamics in a strong field are dominated by the emitted radiation, and a significant fraction of the field energy can be transferred into high-energy radiation.Bulanov et al. (2004); Di Piazza, Hatsagortsyan, and Keitel (2010); Bulanov et al. (2011a); Sokolov, Naumova, and Nees (2011); Ji et al. (2014); Liseykina, Peganov, and Popruzhenko (2025); Gonoskov et al. (2017)
@@
It is rather straightforward to relate the cascade types with the corresponding experimental setups: laser-particle beam interaction is suitable for showers,Pouyez et al. (2024) whereas avalanches can be generated in laser-laser interaction.Bashmakov et al. (2014); Gelfer et al. (2015); Grismayer et al. (2017); Mironov, Gelfer, and Fedotov (2021) It is worth noting that these associations are not set in stone and different types of cascades can occur either almost simultaneously or concurrently in different interaction geometries (e.g., see Refs. Mironov, Narozhny, and Fedotov (2014); Magnusson et al. (2019b); Samsonov, Nerush, and Kostyukov (2019); Samsonov, Kostyukov, and Nerush (2021)).
@@ II.1 Radiation-dominated regime in the classical picture
Let us denote the dimensionless amplitude of a rotating $\mathbf{E}$ field by $a_{*}=2a_{0}$, where
@@
Here, $e>0$ is the elementary charge, $m$ is the electron mass, and $E_{0}$ and $\omega$ are the amplitude and frequency of a single laser beam. Consider an electron injected into the field, initially at rest. According to the classical equations of motion, the electron’s Lorentz factor will oscillate with time as $\gamma_{e}=\sqrt{1+a_{*}^{2}\sin^{2}(\omega t/2)}$. At $a_{*}>1$, the electron will become relativistic within a subcycle of the field. As $a_{*}$ is increased further, the interaction changes qualitatively. Let us briefly discuss these interaction regimes, which we also illustrate in Fig. 1.
@@ II.2 Quantum radiation
where $\mathbf{u}_{e,\gamma}=\mathbf{p}_{e,\gamma}/(mc)$, $\mathbf{p}_{e,\gamma}$ are the electron and photon momenta, and $\mathbf{E}$ and $\mathbf{B}$ are the local electric and magnetic field components of the laser field, respectively. For $e^{\pm}$, $\chi_{e}$ corresponds to the rest frame field strength in units of $E_{S}$. Here, $E_{S}=m^{2}c^{3}/e\hbar$ is the Sauter-Schwinger field. The amplitude $E_{S}$ corresponds to an intensity of about $4.6\times 10^{29}\;\text{W/cm${}^{2}$}$.Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023) We present the full expressions for $W_{\mathrm{nCE}}$ and $W_{\mathrm{nBW}}$ in Appendix A.
@@
Within QED, it is natural to characterize the interaction regimes by the value of $\chi_{e}$. Emission becomes essentially quantum at $\chi_{e}\gtrsim 0.1$, [^2] The radiation intensity calculated within the classical and strong-field QED approaches overlaps at lower values of $\chi_{e}$.Ritus (1985); Baier and Katkov (1968); Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Bulanov et al. (2013); Niel et al. (2018) at which point the particle dynamics change. If the field is well below the Sauter-Schwinger limit ($E\ll E_{S}$) and the electron is ultrarelativistic, the electron motion can still be viewed as semiclassical. Namely, particles propagate along classical trajectories,Baier and Katkov (1968) but the RR is accounted for in random photon emission eventsElkina et al. (2011). These events can be treated as point-like within the locally constant field approximation,Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023) which we assume valid for the interaction regimes discussed in this work, as it generally requires $a_{0}\gg 1$ (see also Appendix B for a more detailed discussion of the approximations in use).
@@
During the electron acceleration, $\chi_{e}$ varies and can increase with time, as well as $\gamma_{e}$. Hence, the emission probability rate is parametrized as $W_{\mathrm{nCE}}(t)=W_{\mathrm{nCE}}[\gamma_{e}(t),\chi_{e}(t)]$, and, more importantly, it can continuously grow along an electron’s semiclassical trajectory. As a result, the field energy can be efficiently converted into radiation. This is a striking feature of the standing-wave field configuration that defines the interaction. Notably, such restoration of $\chi_{e}$ is not revealed, for example, in the interaction of an electron with a single plane wave, as $\chi_{e}$ is conserved during semiclassic motion in between quantum emissions.Landau and Lifshitz The quantum parameter of each particle produced in an elementary QED event (either photon emission or pair production) is smaller than that of the particle initiating the event. These two points combined make the plane-wave setup (namely, a weakly or moderately focused single laser), in principle, unsuited to access the important regimes investigated here. Notably, at very strong focusing, $\chi_{e}$ can grow for electrons injected in the focus.Mironov, Gelfer, and Fedotov (2021) Furthermore, interaction with a dense plasma can change the dynamics significantly, opening access to different interaction regimes, including avalanche-type cascades.Samsonov, Nerush, and Kostyukov (2019); Samsonov, Kostyukov, and Nerush (2021)
Taking this into consideration, we can estimate the time scale $t_{\mathrm{em}}$ at which the accelerated electron emits as $\int^{t_{\mathrm{em}}}W_{\mathrm{nCE}}(t)dt=1$. By $t_{\mathrm{em}}$, the electron acquires $\gamma_{\mathrm{em}}=\gamma_{e}(t_{\mathrm{em}})$ and $\chi_{\mathrm{em}}=\chi_{e}(t_{\mathrm{em}})$. As a result, the properties of emission and the interaction regime are defined by these quantities; $\chi_{\mathrm{em}}$ can be small or large compared to unity, depending on $a_{*}$ and other field parameters. In particular, in a CP field, we can approximate the field dependence at $\chi_{\mathrm{em}}<1$ as:Mercuri-Baron et al. (2025)
@@ II.3 QED cascades
When $\chi_{\mathrm{em}}$ approaches 1 (marked by a star in Fig. 1), the probability of pair creation by the emitted photons becomes high. This signifies the transition from single-electron dynamics in the RD regime to the onset of QED avalanche-type cascades, which is essentially a many-body regime. As the initial and secondary charges continuously experience acceleration and emit hard photons, $e^{-}e^{+}$ pair yield can reach a steady exponential growth in time, $N\sim\exp(\Gamma t)$,Bell and Kirk (2008); Kirk, Bell, and Arka (2009); Fedotov et al. (2010); Mercuri-Baron et al. (2025) where $\Gamma$ is the so-called cascade growth rate. Estimates show that the field scale for reaching high-multiplicity cascades is of the order $a_{*}\sim\alpha a_{S}$Fedotov et al. (2010); Elkina et al. (2011) (shown in Fig. 1). However, it was demonstrated with simulationsRidgers et al. (2012, 2013); Mironov, Narozhny, and Fedotov (2014); Gelfer et al. (2015); Grismayer et al. (2017); Artemenko and Kostyukov (2017); Yu et al. (2018) and recent theoretical analysisMironov, Gelfer, and Fedotov (2021); Mercuri-Baron et al. (2025) that avalanche-type cascades can onset at significantly lower fields.
@@
The spatial distribution of the field, e.g., if laser beams are tightly focused, also affects the cascade growth rate $\Gamma$.Tamburini, Di Piazza, and Keitel (2017) As the interaction area for such fields is small, $e^{\pm}$-s and $\gamma$-s migrate from the focal area during the interaction at rates $\nu_{e,\gamma}\propto\omega$.Jirka et al. (2016) This limits the pair yield in a cascade at lower fields,Tang (2024) and imposes a threshold condition for the steady state formation.Mercuri-Baron et al. (2025) Overall, $\Gamma$ depends on the field parameters parametrically via the rates of particle production and migration:Mercuri-Baron et al. (2025)
@@
$\Gamma\simeq\frac{\omega}{\pi}\left[-1+\frac{1.52\pi^{2}\alpha a_{*}}{\omega\tau_{C}a_{S}}\exp\left(-1.92\frac{\alpha a_{S}}{a_{*}}\right)\right],$ (6)
@@
We emphasise that for efficient particle production, the cascade should reach a steady state. In short laser pulses, this is not guaranteed, especially if the field strength is close to $a_{\mathrm{th}}$. Still, $e^{-}e^{+}$ pair yield can be relatively high already at the onset stage. We denote this subregime as the avalanche precursor: a transition from the quantum RD regime to high-multiplicity avalanches. For the sake of an estimate, one may assume that it extends up to the field strength $a_{*}$ such that $\Gamma(a_{*})=\omega/2\pi$ (shown in Fig. 1 as the vertical green dash-dotted line). This condition guarantees a high particle yield in one field cycle.